OpenStudy (anonymous):

Find all solutions to the equation. (sin x)(cos x) = 0

OpenStudy (anonymous):

@zordoloom Can you help me?

OpenStudy (anonymous):

Sure, would you know how to start?

OpenStudy (anonymous):

No :/

OpenStudy (anonymous):

Alright, so start by splitting it into 2 equations.

OpenStudy (anonymous):

sinx= o cosx=0

OpenStudy (anonymous):

sinx=0 cosx=0 Now solve each equation, would you know how to proceed from here?

OpenStudy (anonymous):

sinx=0 at 0 and pi cosx=0 at pi/2 and 3pi/2

OpenStudy (anonymous):

That's all I know

OpenStudy (anonymous):

Alright, do you know what the unit circle is?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Oh, I didn't see the reply above. So that would be the answer. But, normally it is written a little differently.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

For the first one: x=pi/2+pi(n) and for the second one: x=pi(n)